A transport truck carries stereo, computer, and television parts
the mass of the truck and the parts is 28 t
the mass of the empty truck is 6
the mass of the computer part is 2 t less than the mass of the television parts
the mass of stereo parts is double the mass of television parts
what is the mass of each type of part?
show your work
1
Expert's answer
2020-05-09T14:09:49-0400

truck = 6t

part = stereo + television + computer

computer = televison - 2t

stereo = television * 2


truck + parts = 28t

part = 28t - 6t = 22t


part = computer + television + stereo

part = television - 2t + televison + television*2

part = 4*television - 2t

22t = 4*television - 2t

4*television = 24t


television = 6t

computer = 6t -2t = 4t

stereo = 6t *2 = 12t


television = 6t

computer =4t

stereo = 12t